OpenStudy (anonymous):

a) prove that 1+z+z^2+...+z^n = 1-z^(n+1) / (1-z) for z \(\ne\) 1 b) Using the result of (a), prove that \(1+cos\theta + cos2 \theta + ... + cosn \theta = \frac{1}{2} + \frac{sin(n+1/2)\theta}{2sin(\theta/2)}\)

OpenStudy (anonymous):

I've done part (a), but I don't know how to do part (b). Any ideas?

OpenStudy (loser66):

how do you prove part a) ? by using generating function?

OpenStudy (loser66):

or sum definition?

OpenStudy (anonymous):

^ sum of finite geometric series

OpenStudy (loser66):

the problem ask us use part a to prove part b, it means they must relate to each other, right?

OpenStudy (anonymous):

GS for part (a)

OpenStudy (anonymous):

@RolyPoly: Is the LHS of part b \[1 + \cos \theta + {\cos ^2}\theta + ... + {\cos ^n}\theta \]?

OpenStudy (loser66):

agree with drawar, unclear at that point . verify ,please

OpenStudy (anonymous):

@drawar No, cos(nθ) It is Lagrange trigo's identity.

OpenStudy (anonymous):

So far, I got 1+z+z^2 + ... + z^n = 1-z^(n+1) / (1-z) \[\frac{1-cos(n+1)\theta- i sin(n+1)\theta}{1-(cos+isin\theta)} = 1+cos \theta + cos2\theta + ... + cosn\theta + i(sin\theta + sin2\theta+... + sinn\theta)\]

OpenStudy (anonymous):

That's beyond my knowledge, I betta leave. :)

OpenStudy (loser66):

@drawar do you see my link? exactly this problem

OpenStudy (anonymous):

Yes, but that's has something to do with complex analysis, which I haven't known of...

OpenStudy (loser66):

ok

OpenStudy (anonymous):

\[LS = \frac{1-cos(n+1)\theta- i sin(n+1)\theta}{1-(cos+isin\theta)} \]\[= \frac{1-e^{i(n+1)\theta}}{1-e^{i\theta}} \]

OpenStudy (anonymous):

I didn't know it's complex analysis...

OpenStudy (anonymous):

Anyway, thanks~

OpenStudy (loser66):

you mean you don't understand the way people come up with the problem at that site?

OpenStudy (loser66):

Sorry for that, I just see the problem looks like yours and post the link for you to see. I didn't read it yet. Let see what can I do for you. give me time, I read and IF I can translate in simple way, I will show you mine. is it ok?

OpenStudy (anonymous):

You don't have to say sorry!! Thanks for giving me the link, even though it looks complicated.

OpenStudy (loser66):

hand off, it sounds like only one way to come up with the problem as what people in that link said.

OpenStudy (anonymous):

Maybe... :(

OpenStudy (anonymous):

Thanks again!